What Have We Learned from Studying

High Energy Proton-Antiproton Collisions?

1. The Current Paradigm—The Standard Model 2. The High Energy Frontier

3. What We Have Learned

4. The Next Step

I. Standard Model of Elementary Particles

Particle physics before ∼ 1960 had established

• Elementary particles appeared to be the basic constituents

of universe electron (e±) proton (p) neutron (n)

• Four fundamental interactions or forces:

1. Strong Force—bound nucleons (p, n) together

2. Electromagnetism—bound e− to nuclei

3. Weak Force—responsible for neutron β decay

4. Gravity—most obvious, but also the weakest

• Host of other “elementary particles”

– neutrinos (νe, νμ)

– muon (μ)

– lots of other hadrons (π∘, K±, Δ’s,...) Seemed to be spare wheels....

Over the last 3 decades, many pieces have come together:

• Hadrons are composite objects made up of quarks

• 3 “generations” of quarks and leptons

( \begin{matrix}u \\ d \end{matrix} ) ( \begin{matrix}c \\ s \end{matrix} ) ( \begin{matrix}t \\ b \end{matrix} )  ( \begin{matrix}e \\ νe \end{matrix} ) ( \begin{matrix}μ \\ νμ \end{matrix} ) ( \begin{matrix}τ \\ ντ \end{matrix} )

• Three of the four forces can be described by relativistic quantum field theories:

– “bosons” mediate each force

EM  γ
Weak  W + , W −, Z∘
Strong  gluons

Breakthrough was Electro-Weak Unification

• A single theory that “unified” the EM and Weak force – γ, W ± and Z∘ are part of one “gauge” group

– W ± and Z∘ bosons mediate the weak force

– their masses come from interaction with a Higgs particle

– photon remains massless

⇒ the “Standard Model” or Weinberg-Salam-Glashow Theory

Successes and Failures

Model is very successful in describing Electroweak data

• A large body of data has been gathered to test model – Low- and high-energy νe scattering

– Phenomenology of weak decays

– Properties of the gauge bosons

– Electromagnetic-Weak interference effects

No hard data that is in fundamental disagreement with the SM

So what’s wrong?

1. Too many numbers!!

Model contains an inordinate number of parameters: • 11 fermion and boson masses

• 3 gauge couplings

• 4 quark mixing angles

2. Leaves too many questions unanswered: • What generates the fermion mass scales?

• Why are there 3 generations?

• Is there a unified description of all forces?

3. Incomplete for very high energy interactions

• Weak force becomes “strong” at an energy scale of O(1000) GeV

• The “Heirarchy” problem

Options for Making Progress

1. Wait for a bright idea

• String theory (theory of “everything”)

• Supersymmetry

• Technicolor

2. Test assumptions and predictions of the model

• make measurements that test prediction, e.g. For example,

sin2 θW ≡ 1 − (\frac{MW}{MZ} )2  (1)
= \frac{πα}{√2GμMW 2}[\frac{1}{1 − Δr(Mtop, MH)} ] (2)

where:

– MW, MZ are W and Z boson masses – Gμ is μ decay constant

– Δr is “radiative” correction

3. Find all the predicted particles

(a) the top quark

(b) the Higgs particle(s)

4. Search for new phenomena

• structure of quarks and leptons • new forces

• new classes of particles

Clearly would like to pursue all of these...

So Why High Energy Collisions?

Most searches/measurements require very energetic initial states

For example,

• Only practical way of producing large numbers of W bosons is via quark-antiquark collisions:

• A new force could possible be mediated by very heavy bosons:

• Structure inside a quark would show up when quarks interact at very high energies

The Highest Energy Colliders

Fermilab Tevatron Collider  p p  √s = 1800 GeV
CERN sp ps  p p  √s = 630 GeV
LEP  e+ e−  √s = 100 GeV
SLAC SLC  e+ e−  √s = 100 GeV

The Fermilab Collider is on the energy frontier:

• Collider has 6 bunches of protons and anti-protons circulating in opposite directions

⇒ p p interaction rate of ∼ 50 kHz

• Large body of data recorded by a detector called CDF – ∼ 5 × 106 events recorded in 1988-89

– Analysis of data nearing completion

The CDF Detector

• built by a collaboration of ∼ 250 physicists • general-purpose, 5000 ton detector

– Calorimeter to measure the energies of the particles – A magnetic spectrometer that measures charged

particle momenta

– A sophisticated real-time triggering system that records on average about 1 in every 50,000 collisions

The most interesting collisions are those that generate high-p_{T} particles

• These are filtered from the background events by requiring one or more energetic objects in the event

– An e± or μ± with pT > 12 GeV

– A neutrino with ET > 25 GeV

– Jet with ET > 120 GeV

Experimental Complications

Transverse Variables (E_{T}, etc.)

• Initial state is a collision between a parton inside one proton and a parton inside one anti-proton

• The density of particles in p p collisions rises like

∼ ln θ  (3)

but distribution in pT remains approximately constant

Physics Results from CDF

Experiment is really several quasi-independent studies:

1. Electroweak Physics (W ’s and Z’s)

2. Heavy Flavour Physics (top and bottom quark studies)

3. Non Standard Model Studies (SUSY, W ′’s, etc)

4. Perturbative QCD (jets)

5. Non-perturbative QCD (minimum-bias)

Will focus on work that is pushing the frontier

• Electroweak studies using the – ∼ 5000 W → eνe and W → μνμ events – ∼ 400 Z → e+ e− and Z → μ+ μ− events

• New particle searches (top, W ′’s, etc.)

Electroweak Physics

Properties of the W and Z bosons:

• Measured the mass and width of the Z

– used subsample of the best measured electrons and muons MZ = 91.9 ± 0.3(stat. + syst ) ± 0.2(scale ) GeV/c2

• Measured the mass and width of the W

W → e^{±}ν_{e} and W → μ^{±}ν_{μ}  (4)

– Infer MW from the “transverse mass”

M{eν \atop T} ≡ √2ET E̸T (1 − cos ϕeν)

Use of E̸T introduces systematic uncertainties associated with understanding calorimeter response

– 1130 eνe and 592 μνμ candidates

MW = 79.91 ± 0.39 GeV/c2

Results in a measurement of sin2 θW = 0.231 ± 0.008

Other Electroweak studies have been completed:

• Cross section ratio

\frac{σ(W → eνe)}{σ(Z → ee)} = \frac{σ(pp → W X)}{σ(pp → Z)} \frac{Γ(W → eν)}{Γ(Z → ee)} \frac{Γ(W)}{Γ(Z)}

gives another test of the standard model

– \frac{σ(W →eνe)}{σ(Z→ee)} = 10.2 ± 0.8 ± 0.4

– Using LEP Γ(Z) measurement, can determine Γ(W) = 2.19 ± 0.20 GeV

– If assume Nν = 3, then Mtop> 43 GeV/c2 at 95% CL.

• Z → e+ e− asymmetry – measure of the left and right-handed gauge couplings

• W → e+ νe asymmetry – indirect measure of the parton distribution functions

• Direct σ(W) and σ(Z) measurements – tests QCD and Standard Model

The Top Quark Search

Existence of top is a fundamental SM prediction

• riddled with inconsistencies if a new quark doesn’t exist with a mass < ∼ 180 − 200 GeV/c2

• Tevatron is the only place where it can be found

Search for events with unusual properties

• Inclusive lepton final states

• Dilepton final states

• Combined multi-channel analysis

Search in Inclusive Lepton Channel

We searched for top by requiring

• A “golden” electron candidate

• Require E̸T > 20 GeV to tag presence of energetic neutrino • At least two jets with |η| < 2.0 and ET > 10 GeV

⇒ 104 candidate events

Studied the backgrounds in detail

• Removed conversions using tracking information • Rate of QCD fakes and interacting π+ < ∼ 10% • Events from semi-leptonic b decays < ∼ 10%

• Primary background comes from real pp → W + jets

Removed W + jets background statistically by fitting transverse mass distribution to

\frac{dN}{dM_{T}^{eν}} = αT (M_{T}^{eν}) + βW (M_{T}^{eν}).

To verify our understanding of M_{T}^{eν}, we also studied the W + 1 jet data

• t t contribution small because of 1 jet requirement

• Good agreement confirms that we understand M{eν \atop T}

The dilepton search in the eμ channel found one candidate event

• It is a very clean event

• also has a forward μ candidate (from B?)

• not typical of primary background (Z → τ+ τ−)

• Assuming candidate is real t t, set mass limit of 28 < Mtop < 72 GeV/c2 (at 95% CL)

Subsequently extended limit to exclude up to 91 GeV/c^{2}

• Used e+ e− and μ+ μ− after removing Z’s

• Used inclusive lepton + jets by requiring a low-energy μ candidate from B decay

Exotic Particle Searches

Searched for exotic phenomena:

• Extensions of standard model (W ′’s and Z′’s) – Search for analomolous high mass events

– See no excess of events in lepton + neutrino channels – Set limit of

MW ′ > 520 GeV/c2 (95%C.L.)

– See no evidence in either e+ e− or μ+ μ− channels – set limit of

MZ′ > 412 GeV/c2 (95%C.L.)

Also have looked at the most energetic events

• Attempt to see if there is anything “unusal”

• However, appear to be consistent with QCD predictions

A Look into the Future

CDF will acquire about 25 times more data in next two years

• W mass measurement with σ ∼ 100 MeV/c2

• Top search should reach a mass as large as 150-160 GeV/c2 • Increase sensitivity to other “new” physics

Will still not address several open questions:

• The existence of the Higgs boson

• What happens at O(1) TeV?

The Multi-Tev Energy Scale

These questions can be answered at the SSC

• Provide us access to the multi-TeV energy scale

• A copious source of top quarks

• Push our search for new phenomena

Several experiments being proposed

• Will probably be two large general-purpose detectors – SDC – GEM

• Several more specialized experiments later

So What Have We Learned?

The Standard Model is alive and well

• W boson mass and width

• W and Z production and decay asymmetries

• QCD describes high PT jet production

Have to look harder for New Physics

• Top not yet in sight

• No evidence for new interactions

Have yet to reach the bottom of the barrel...

• Can improve precision measurements

• New particle searches limited only by running time